A gas is compressed at a constant pressure of 0.800 atm from 10.00 L to 1.00 L. In the process, 340 J of energy leaves the gas by heat.

(a) What is the work done on the gas?
Answer #1

Work done by the gas is given by PdV i.e. pressure x change in volume

1 Litre = 10^-3 m^3 hence

work done = 0.8 x (1-10) x 10^-3 = 7.2 x 10^-3 atm m^3

now 1 atm = 10^5 N/m^2 =

hence work done = 720 J